The effective cross sectional area of a horse's femur is 7×10^-4m^2 and the Young's modulus of this bone is 8.3×10^9pa.Calculate the strain that occurs in the femur when the horse of mass 600kg puts it's full weights on it
1
Expert's answer
2019-07-10T13:31:18-0400

The Hooke's law states

"\sigma=\varepsilon E"

The stress is given by relation

"\sigma=\frac{F}{A}"

So, strain

"\varepsilon=\frac{\sigma}{E}=\frac{F}{AE}=\frac{mg}{AE}"

"=\frac{600\:\rm{kg}\times 9.8\:\rm{N/kg}}{7\times 10^{-4}\:\rm{m^2}\times 8.3\times 10^9\:\rm{Pa}}"

"=0.001\:\rm{m}=1\:\rm{mm}"